# Sharp Eigenfunction Bounds on the Torus for large $p$

### 11 Mar 2026 v1

@@ Abstract.
We prove the discrete restriction conjecture holds with no loss when $p>\frac{2d}{d-4}$ and $d\geq 5$. That is, we show optimal $L^{p}$ bounds for eigenfunctions of the Laplacian on the square torus for large values of $p$. This improves the results Bourgain and Demeter. Our proof method is a refinement of the circle method approach previously used to establish results with a subpolynomial loss. This represents the first sharp $L^{p}$ bounds for eigenfunctions on the torus since the work of Cooke and Zygmund.
We present applications to bounds for spectral projectors and the additive energy of lattice points on higher dimensional spheres. These results are similarly sharp. We also prove results with a logarithmic loss that hold in a wider range of $p$.
@@ 1.1. Background
Let $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ be the square torus. We may view functions on this manifold as $\mathbb{Z}^{d}$-periodic functions on $\mathbb{R}^{d}$, whose values are completely determined by its values on a fundamental domain $0,1]^{d}$. If $f\in L^{2}(\mathbb{T}^{d})$, it is well known we may expand this function in Fourier series.
@@
The plane waves $e^{2\pi i\mathbf{k}\cdot x}\coloneq e(\mathbf{k}\cdot x)$ are eigenfunctions [^1] We use $\mathbf{k}$ to represent integer vectors in $\mathbb{Z}^{d}$. For other notional conventions, see Section 1.5 of the square root of the Laplacian $\sqrt{-\Delta}$ with eigenvalue $2\pi|\mathbf{k}|$. It would be equivalent in what follows to consider eigenfunctions of $\Delta$, although statements would have altered numerology. Let $P_{N}$ be the projection onto the eigenspace of $\sqrt{-\Delta}$ with eigenvalue $2\pi N$. Then
(2) $e_{N}(x)=P_{N}f(x)=\sum_{|\mathbf{k}|=N}\mathcal{F}(f)(\mathbf{k})e^{2\pi i\mathbf{k}\cdot x}.$
That is, we project onto integer lattice point frequencies that lie on the circle of radius $N$. We will also use the notation $\lambda=N^{2}$. Bourgain in [2] asked how large such functions can be in $L^{p}$.
@@
Let $p\geq 2$ and $d\geq 2$. Let $\mathbb{T}^{d}$ be the square torus. Suppose $e_{N}$ is an eigenfunction of $\sqrt{-\Delta}$. Then
@@
The extra $O_{\epsilon}(N^{\epsilon})$ factor can be omitted if $d\geq 5$.
@@
Then it is well known that the following bounds hold for the cardinality of lattice points.
@@
Conjecture 1.1 was proven with a factor of $C_{\epsilon}N^{\epsilon}$ in all dimensions when $p\leq\frac{2(d+1)}{d-1}$ by the $\ell^{2}$-decoupling theorem of Bourgain and Demeter [5]. It was also proven for $p\geq\frac{2d}{d-3}$ by the same authors in [4], also with an $\epsilon$-loss. Combining the results of these two papers yields the conjecture with loss when $p\geq\frac{2(d-1)}{d-3}$. Our main result is that this loss can be removed for large $p$.
@@
This result contains the first sharp bounds for toral eigenfunctions since the work of Cooke and Zygmund. Cooke proved the result in the course of showing a Cantor-Lebesgue Theorem [7]. The result was studied in its own right in [15] to answer a question of Fefferman. Both proved the sharp result with no loss when $d=2$ and $2\leq p\leq 4$ using essentially the same geometric techniques. Our proof method differs greatly from that work. In particular, we refine the circle method approach used in [2] and [4]. For this reason, our results are limited to the square torus. As an aside, other works often refer to bounds with a power of $N^{\epsilon}$ as “sharp”. In this paper, we will reserve sharp to mean that no such factor is present although, for low dimensions and large $p$, the factor is necessary.
We also able to improve the $N^{\epsilon}$ loss to a logarithmic loss in a larger range of $p$, but only for $d\geq 6$.
@@
Let $d\geq 6$ and $p>\frac{2d}{d-3}$ or $d=5$ and $p>6$. Let $\mathbb{T}^{d}$ be the square torus. Then for every eigenfunction $e_{N}$, the following bound holds.
@@ 1.2. Organization of the paper
In Section 1.3 we present some immediate corollaries of our main theorem which may be of independent interest. In Section 1.4 we give a tour of the discrete restriction conjecture that intuitively establishes the numerology of the conjecture and when an $N^{\epsilon}$ factor is necessary. Section 1.5 contains the notation we use.
In Section 2 we present the proof of Theorem 1.3 and in Section 3 we present the proof of Theorem 1.2. These sections are written to be largely independent of each other. Throughout both proofs we use certain known exponential sum estimates. We sketch their proofs in Section 4. As our proof is broadly a refinement of the work of Bourgain and Demeter, specifically in [4], we take Section 5 to list the differences between our proof and theirs and to note the limitations that prevent us from extending the sharp result to the full range of $p$ for which the result with $\epsilon$-loss is known.
@@ 1.3. Some Applications
We present two applications of our main results to bounds for approximate eigenfunctions and bounds for the additive energy of spheres.
Instead of studying $L^{2}\rightarrow L^{p}$ bounds for projections onto eigenspaces, one could instead consider projections onto spectral bands. This is equivalent to estimating the $L^{p}$ norm of $L^{2}$-normalized quasimodes. For $1\leq\delta\lesssim\lambda^{-1}$ define
@@
Here $A_{N,\delta}$ is an annulus of radius $\sim N$ and width $\delta$. There is no benefit to projecting onto smaller windows as when $\delta\leq\frac{1}{4}\lambda^{-1}$; this recovers $P_{N}$ defined in (2). Germain and Myerson asked for the sharp $L^{2}\rightarrow L^{p}$ bounds for this object in [10], after the problem was first studied by Hickman [11]. Sharp estimates for $p\leq\frac{2(d+1)}{d-1}$ for $\delta$ away from the eigenfunction case were proven in [13] for all tori, not just square ones.
We confirm the sharp version of their conjecture in the analogous range.
@@
Let $d\geq 5$ and $\mathbb{T}^{d}$ be the square torus. Let $p>\frac{2d}{d-4}$ and let $P_{N,\delta}$ be the spectral projector with $1\leq\delta\lesssim\lambda^{-1}$. Then
@@
Additive energies capture the additive structure of a set. Approximate groups maximize the additive energy for sets of a fixed size, for example. For $d\geq 5$, we give improved results for the additive energy of spheres.
Before stating our results, we quickly derive the conjectured bound. For $d\geq 5$, we have $\mathcal{F}_{d,N}\sim N^{d-2}$. Consider $n\mathcal{F}_{d,N}$, the set formed by adding $n$ elements of $\mathcal{F}_{d,N}$. Each element in this set has magnitude $\lesssim_{m}N$, and $n\mathcal{F}_{d,N}\lesssim N^{n(d-2)}$. If the sumsets equidistributed over lattice points in a ball of radius $\lesssim_{m}N$, each point would have $\sim N^{n(d-2)-d}$ $m$-tuples that sum to that given point. An application of Cauchy-Schwarz gives for $d\geq 5$ the lower bound
@@
We have $E_{n}(\mathcal{F}_{N,d})\lesssim N^{2n(d-2)-d}$ when
@@
Additionally, the upper bound holds with a $(\log N)^{\frac{d-2}{d-4}}$ factor if
@@
Here, $K(x)$ is the convolution kernel of $P_{N}$, which is given by (7) in the case $\mathcal{F}(f)(\mathbf{k})=1$ for every $\mathbf{k}$. The result then follows from an application of Theorem 1.2 or Theorem 1.3 depending on which is applicable in the given range of $p$ and $d$.
@@ 1.4. A quick tour of Discrete Restriction
Clearly $f(0)=M$. By continuity, we have $|f(0)|\geq M/2$ is a ball of radius $\sim N^{-1}$ centered at 0. We then have
@@
The constructive interference of the exponentials, which is a manifestation of the uncertainty principle, forces a function that is very large on a small neighborhood of a point. In the case of the torus, when $d\geq 5$, we have $M\sim N^{d-2}$. The same computation then gives the first bound on the right hand side of Conjecture 1.1. When $d=2,3$ or $4$, we have a factor of $N^{\epsilon}$.
This point focusing example is a general phenomena that saturates $L^{p}$ norms of eigenfunctions (or, more generally, quasimodes). However, the two terms in the discrete restriction conjecture balance at $p=\frac{2d}{d-2}$. On generic manifolds we expect the critical value to be $p=\frac{2(d+1)}{d-1}$. The difference is due to the degenerate number of lattice points on spheres as compared to arbitrary $1$-separated sets, which gives a global count of $N^{d-2}$ as opposed to $N^{d-1}$.
The second term on the right hand side of Conjecture 1.1 can be seen by considering a single plane wave. For an example with full support, one can generate a host of random examples analyzed with Khintchine’s inequality that experience maximum destructive interference. The difference in the saturating example leads to the differences in techniques used for $p>\frac{2d}{d-2}$ and $p<\frac{2d}{d-2}$. Harmonic analysis techniques seem most suited for smaller value of $p$, while number theory works well with the analysis of exponential sums experiencing constructive interference. See [9] for an exploration of this dichotomy in the case of spectral projectors. There, sharp estimates are proven both using the circle method and cancellation between exponential sums.
@@ 1.5. Notation
We denote $\mathbf{k}$ a vector in $\mathbb{Z}^{d}$ with components $(k_{1},...,k_{d})$. The variable $x$ will denote a point in $\mathbb{T}^{d}$. We often abuse notation to identify this with a point in the fundamental domain $0,1]^{d}$ with opposite sides identified. This is most commonly done in the $d=1$ case to identify $[0,1]$ and $S^{1}$.
@@
We will write $e(y)=e^{2\pi iy}$. For a function defined on $\mathbb{R}^{d}$ we will write $\hat{f}$ for the Fourier transform:
@@ 1.6. Thanks
The author would like to thank Xiaoqi Huang and Connor Quinn for careful readings of this and earlier draft of this manuscript.
@@ 2.1. Preliminary estimates
The constant $c_{Q}$ is chosen so that $\int_{[0,1]}\eta_{Q}dt=1$. Note that $c_{Q}\sim\log Q$ by direct computation. This will help us get an improved bound for Fourier coefficients on our error term later. We now define
@@
Here, $G(t,x_{j})=\sum_{k\in\mathbb{Z}}\gamma(k/N)e(kx_{j}+k^{2}t)$ is the 1-dimensional Schrödinger propagator on the square torus. We now seek to bound the size of $K^{Q}$ and the size of the Fourier transform of $K-K^{Q}$. To achieve the first task, we need a bound on $G(t,x)$ (we drop the subscript $i$ momentarily). A sharp bound can be proven for all time using classical methods, however we also require an estimate that applies near fractions of very large denominator.
@@
Let $q\geq N$ be an odd integer. Let $t=\frac{a}{q}+\varphi$ where $(a,q)=1$ and $|\varphi|\lesssim\frac{1}{q^{2}}$ Then
(13) $|G(t,x)|\lesssim\sqrt{q}.$
@@
Because of this we may integrate by parts. Note
$J(x,\varphi,m,q)=\int_{\mathbb{R}}\gamma(y/N)\frac{1}{L^{\prime}(y)}\frac{d}{dy}e(L(y))dy.$
@@
$J(x,\varphi,m,q)=-\int_{\mathbb{R}}\Bigl(\frac{\gamma^{\prime}(y/N)}{NL^{\prime}(y)}-\frac{\gamma(y/N)L^{\prime\prime}(y)}{(L^{\prime}(y))^{2}}\Bigr)e(L(y))dy.$
@@
We now state two related lemmas that work for $q\geq N$.
@@
Let $N\leq Q\leq N^{2}$ and $K^{Q}$ be as before. Then for $d\geq 4$, the following $\log$ loss version of Proposition 4.2 in [4] holds.
(16) ${K^{Q}}||_{\infty}\lesssim\log QN^{2}Q^{\frac{d-4}{2}}.$
@@
Observing (15), if $|G(t,x)|\gtrsim 2^{s}$ we immediately have
@@
as we have localized to small neighborhoods are rationals with large denominators. We then immediately have via Hölder’s inequality and our normalization constant $c_{Q}$:
@@
After exchanging the order of integrating, we see that the integral with respect to $x$ is only non-zero when $\mathbf{k}=(k_{1},...,k_{d})$ where $k_{j}$ come from the sum in the definition of $G(t,x_{j})$. This then leaves us with only a Fourier transform with respect to time, and so

### 06 Apr 2026 v2

@@ Abstract.
We prove the discrete restriction conjecture holds with no loss when $p>\frac{2d}{d-4}$ and $d\geq 5$. That is, we show optimal $L^{p}$ bounds for eigenfunctions of the Laplacian on the square torus for large values of $p$. This improves the results of Bourgain and Demeter. Our proof method is a refinement of the circle method approach previously used to establish results with a subpolynomial loss. This represents the first sharp $L^{p}$ bounds for eigenfunctions on the torus since the work of Cooke and Zygmund.
We present applications to bounds for spectral projectors and the additive energy of integer lattice points on higher dimensional spheres. These results are similarly sharp. We also prove results with a logarithmic loss that hold in a wider range of $p$.
@@ 1.1. Background
Let $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ be the square torus. We may view functions on this manifold as $\mathbb{Z}^{d}$-periodic functions on $\mathbb{R}^{d}$, whose values are completely determined by their values on the fundamental domain $0,1]^{d}$. If $f\in L^{2}(\mathbb{T}^{d})$, it is well known we may expand this function in Fourier series.
@@
The plane waves $e^{2\pi i\mathbf{k}\cdot x}\coloneq e(\mathbf{k}\cdot x)$ are eigenfunctions [^1] We use $\mathbf{k}$ to represent integer vectors in $\mathbb{Z}^{d}$. For other notional conventions, see Section 1.5 of the square root of the negative Laplacian $\sqrt{-\Delta}$ with eigenvalue $2\pi|\mathbf{k}|$. It would be equivalent in what follows to consider eigenfunctions of $\Delta$, although statements would have altered numerology. Let $P_{N}$ be the projection onto the eigenspace of $\sqrt{-\Delta}$ with eigenvalue $2\pi N$. Then
(2) $e_{N}(x)\coloneq P_{N}f(x)=\sum_{|\mathbf{k}|=N}\mathcal{F}(f)(\mathbf{k})e^{2\pi i\mathbf{k}\cdot x}.$
That is, we project onto integer lattice point frequencies that lie on the sphere of radius $N$. We will also use the notation $\lambda=N^{2}$. Bourgain in [2] asked how large such functions can be in $L^{p}$.
@@
Let $p\geq 2$ and $d\geq 2$. Let $\mathbb{T}^{d}$ be the square torus. Suppose $e_{N}$ is an eigenfunction of $\sqrt{-\Delta}$. Let $\epsilon>0$. Then
@@
The extra $C_{\epsilon}N^{\epsilon}$ factor can be omitted if $d\geq 5$.
@@
It is well known that the following bound hold for the cardinality of integer lattice points on the sphere.
@@
Conjecture 1.1 was proven with a factor of $C_{\epsilon}N^{\epsilon}$ in all dimensions when $p\leq\frac{2(d+1)}{d-1}$ by the $\ell^{2}$-decoupling theorem of Bourgain and Demeter [5]. It had been previously proven for $p\geq\frac{2d}{d-3}$ by the same authors in [4], also with an $\epsilon$-loss. Combining the results of these two papers yields the conjecture with loss when $p\geq\frac{2(d-1)}{d-3}$. Our main result is that this loss can be removed for large $p$.
@@
This result contains the first sharp bounds for toral eigenfunctions since the work of Cooke and Zygmund. Cooke proved the result in the course of showing a Cantor-Lebesgue Theorem in [7]. The result was studied on its own by Zygmund in [15] to answer a question of Fefferman. Both proved the sharp result with no loss when $d=2$ and $2\leq p\leq 4$ using essentially the same geometric argument. Our proof method differs greatly from that work. In particular, we refine the circle method approach used in [2] and [4]. For this reason, our results are limited to the square torus. As an aside, other works often refer to bounds with a power of $N^{\epsilon}$ as “sharp”. In this paper, we will reserve sharp to mean that no such factor is present, although, for low dimensions and large $p$, the factor is necessary.
We also able to improve the $N^{\epsilon}$ loss to a logarithmic loss in a larger range of $p$.
@@
Let $d\geq 6$ and $p>\frac{2d}{d-3}$ or $d=5$ and $p>6$. Let $\mathbb{T}^{d}$ be the square torus. Then, for every eigenfunction $e_{N}$, the following bound holds.
@@ 1.2. Organization of the paper
In Section 1.3 we present some immediate corollaries of our main theorem which may be of independent interest. In Section 1.4 we give a tour of the discrete restriction conjecture that intuitively establishes the numerology of the conjecture and when a $C_{\epsilon}N^{\epsilon}$ factor is necessary. Section 1.5 contains the notation we use.
In Section 2 we present the proof of Theorem 1.3 and in Section 3 we present the proof of Theorem 1.2. These sections are written to be largely independent of each other. Throughout both proofs we use certain known exponential sum estimates. We sketch their proofs in Section 4. As our proof is broadly a refinement of the work of Bourgain and Demeter, specifically [4], we take Section 5 to list the differences between this proof and theirs.
@@ 1.3. Some Applications
We present two applications of our main results: bounds for approximate eigenfunctions and bounds for the additive energy of spheres.
Instead of studying $L^{2}\rightarrow L^{p}$ bounds for projections onto eigenspaces, one could instead consider projections onto spectral bands. This is equivalent to estimating the $L^{p}$ norm of $L^{2}$-normalized quasimodes. For $1\geq\delta\gtrsim\lambda^{-1}$, define
@@
Here $A_{N,\delta}$ is an annulus of radius $N$ and width $\delta$. There is no benefit to projecting onto smaller windows as, when $\delta\leq\frac{1}{4}\lambda^{-1}$, the projector equals $P_{N}$ defined in (2). Germain and Myerson asked for the sharp $L^{2}\rightarrow L^{p}$ bounds for this object in [10], after the problem was first studied by Hickman [11]. Sharp estimates for $p\leq\frac{2(d+1)}{d-1}$ for $\delta$ away from the eigenfunction case were proven in [13] for all tori, not just square ones.
We confirm the sharp version of the conjecture of Germain and Myerson in the analogous range.
@@
Let $d\geq 5$ and $\mathbb{T}^{d}$ be the square torus. Let $p>\frac{2d}{d-4}$ and let $P_{N,\delta}$ be the spectral projector with $1\geq\delta\gtrsim\lambda^{-1}$. Then
@@
Additive energies capture the additive structure of a set. Approximate groups maximize the additive energy for sets of a fixed size. For $d\geq 5$, we give sharp results for the additive energy of spheres.
Before stating our results, we quickly derive the conjectured bound. For $d\geq 5$, we have $\mathcal{F}_{d,N}\sim N^{d-2}$. Consider $n\mathcal{F}_{d,N}$, the set formed by adding $n$ elements of $\mathcal{F}_{d,N}$. Each element in this set has magnitude $\lesssim_{n}N$, and $n\mathcal{F}_{d,N}\lesssim N^{n(d-2)}$. If the sumsets equidistributed over lattice points in a ball of radius $\lesssim_{n}N$, each point would have $\sim N^{n(d-2)-d}$ $n$-tuples that sum to that given point. An application of Cauchy-Schwarz gives, for $d\geq 5$, the lower bound
@@
We have $E_{n}(\mathcal{F}_{N,d})\sim N^{2n(d-2)-d}$ when
@@
Additionally, an upper bound holds with a $(\log N)^{\frac{d-2}{d-4}}$ factor if
@@
Here, $K(x)$ is the convolution kernel of $P_{N}$, which is given by (7) in the case $\mathcal{F}(f)(\mathbf{k})=1$ for every $\mathbf{k}$. The result then follows from an application of Theorem 1.2 or Theorem 1.3, depending on which is applicable in the given range of $p$ and $d$.
@@ 1.4. A quick tour of Discrete Restriction
Clearly, $f(0)=M$. By continuity, we have $|f(0)|\geq M/2$ in a ball of radius $\sim N^{-1}$ centered at 0. We then have
@@
The constructive interference of the exponentials, which is a manifestation of the uncertainty principle, forces a function that is very large on a small neighborhood of a point. In the case of the torus, when $d\geq 5$, we have $M\sim N^{d-2}$. The same computation then gives the first bound on the right hand side of Conjecture 1.1. When $d=2,3$ or $4$, we have a factor of $C_{\epsilon}N^{\epsilon}$.
This point focusing example is a general phenomena that saturates $L^{p}$ norms of eigenfunctions (or, more generally, quasimodes). However, the two terms in the discrete restriction conjecture balance at $p=\frac{2d}{d-2}$. On generic manifolds, we expect the critical value to be $p=\frac{2(d+1)}{d-1}$. The difference is due to the degenerate number of lattice points on spheres as compared to arbitrary $1$-separated sets, which gives a global count of $N^{d-2}$ as opposed to $N^{d-1}$.
The second term on the right hand side of Conjecture 1.1 can be seen by considering a single plane wave. For an example with full support, one can generate a host of random examples analyzed with Khintchine’s inequality that experience maximum destructive interference. The differences in the saturating examples lead to the differences in techniques used for $p>\frac{2d}{d-2}$ and $p<\frac{2d}{d-2}$. Harmonic analysis techniques seem most suited for smaller value of $p$, while number theory works well with the analysis of exponential sums experiencing constructive interference. See [9] for an exploration of this dichotomy in the case of spectral projectors. There, sharp estimates are proven both using the circle method and cancellation between exponential sums.
@@ 1.5. Notation
We denote $\mathbf{k}$ a vector in $\mathbb{Z}^{d}$ with components $(k_{1},...,k_{d})$. The variable $x$ will denote a point in $\mathbb{T}^{d}$. We often abuse notation to identify this with a point in the fundamental domain $0,1]^{d}$. This is most commonly done in the $d=1$ case to identify $[0,1]$ and $S^{1}$.
@@
We will write $e(y)=e^{2\pi iy}$. For a function defined on $\mathbb{R}^{d},$ we will write $\hat{f}$ for the Fourier transform:
@@ 1.6. Thanks
The author would like to thank Xiaoqi Huang and Connor Quinn for careful readings of this and earlier drafts of this manuscript. The author would also like to thank Simon Myerson for clarifying conversations on exponential sums which improved the presentation of Section 4, and Jonathan Hickman for pointing out an error in the definition of the local operator in Section 3.
@@ 2.1. Preliminary estimates
The constant $c_{Q}$ is chosen so that $\int_{[0,1]}\eta_{Q}dt=1$. Note that $c_{Q}\sim\log Q$ by direct computation. This will help get an improved bound for Fourier coefficients on the error term later. We now define
@@
Here, $G(t,x_{j})=\sum_{k\in\mathbb{Z}}\gamma(k/N)e(kx_{j}+k^{2}t)$ is the 1-dimensional Schrödinger propagator on the square torus. We now seek to bound the size of $K^{Q}$ and the size of the Fourier transform of $K-K^{Q}$. To achieve the first task, we need a bound on $G(t,x)$ (we drop the subscript $i$ momentarily). A sharp bound can be proven for all time using classical methods, however, we also require an estimate that applies near fractions of very large denominator.
@@
Let $q\geq N$ be an odd integer. Let $t=\frac{a}{q}+\varphi$ where $(a,q)=1$ and $|\varphi|\lesssim\frac{1}{q^{2}}$. Then
(13) $|G(t,x)|\lesssim q^{1/2}.$
@@
Because of this, we may integrate by parts. Note
$J(x,\varphi,m,q)=\int_{\mathbb{R}}\gamma(y/N)\frac{1}{L^{\prime}(y)}\frac{d}{dy}e^{L(y)}dy.$
@@
$J(x,\varphi,m,q)=-\int_{\mathbb{R}}\Bigl(\frac{\gamma^{\prime}(y/N)}{NL^{\prime}(y)}-\frac{\gamma(y/N)L^{\prime\prime}(y)}{(L^{\prime}(y))^{2}}\Bigr)e^{L(y)}dy.$
@@
We now state two related lemmas that work for $q\leq N$.
@@
Let $N\leq Q\leq N^{2}$ and $K^{Q}$ be as before. Then, for $d\geq 4$, the following $\log$ loss version of Proposition 4.2 in [4] holds.
(16) ${K^{Q}}||_{\infty}\lesssim\frac{\log Q}{\log(NQ^{-1/2})}N^{2}Q^{\frac{d-4}{2}}.$
@@
Observing (15), if $|G(t,x)|\gtrsim 2^{s}$, we immediately have
@@
as we have localized to small neighborhoods are rationals with large denominators. We then immediately have via Hölder’s inequality and our normalization constant $c_{Q}$ that
@@
After exchanging the order of integration, we see that the integral with respect to $x$ is only non-zero when $\mathbf{k}=(k_{1},...,k_{d})$ where $k_{j}$ come from the sum in the definition of $G(t,x_{j})$. This then leaves us with only a Fourier transform with respect to time, and so